# Transitive Anosov flows on non-compact manifolds

### v1

@@ Abstract.
In this article we study topological transitivity of Anosov flows on non-compact 3-manifolds. We provide homological conditions under which the lifts of a transitive Anosov flow to certain infinite covers of the manifold remain transitive. With some deep results in 3-manifold topology, we then deduce that such cover can be obtained for any non-graph manifold admitting a transitive Anosov flow. Moreover, for a large class of Anosov flows known as $\mathbb{R}$-covered, which are always transitive, we show that their lifts to any regular covers are either transitive or consist exclusively of wandering orbits. Finally, we construct a family of transitive Anosov flows on non-compact manifolds that satisfy a homotopical characterization of suspension flows, answering the flow version of a question concerning the existence of transitive Anosov diffeomorphisms on non-compact manifolds.
@@ 1. Introduction
The topological study of Anosov flows, or Anosov diffeomorphisms, has been so far almost exclusively limited to the case where the ambient manifold is compact. Some exceptions to that rule can be found in the works on Anosov diffeomorphisms on the plane or open surfaces, e.g., [GN15, Mat21, Men77, OD24], as well as recently, in the introduction and study of Anosov-like actions [BFM25, BBM24], where compactness is not explicitly needed.
@@
One of the issues appearing when considering Anosov dynamics in the non-compact setting is that the non-wandering set, i.e., the part of the manifold where the interesting dynamics appear, can be small or even trivial. (A classical example is that there exists a complete metric making the translation on $\mathbb{R}^{2}$ Anosov [Whi73].) So for the dynamics to be interesting, one would like to be able to ensure that the non-wandering set is large enough. In this article, we consider (regular) infinite covers of compact manifolds, as they are a very large and natural source of examples of Anosov flows in the non-compact setting and address the following general question:
@@
Let $\varphi$ be a transitive Anosov flow on a compact $3$-manifold $M$. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be any (infinite) cover of $M$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ the lifted flow. How common is it for $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ to be transitive?
This question of existence and/or abundance of transitive Anosov flows in non-compact manifolds, or infinite covers of compact manifolds has appeared more or less explicitly in several different works: In [DS21], they consider certain Abelian covers $Y$ of an Anosov flow $\varphi$ on $M$ and assume the transitivity of the lifted flow $\varphi_{Y}$, making it natural to wonder whether or not this assumption is necessary [^1] Contrarily to this article, the dimension of the manifolds they consider in [DS21] is arbitrary.. In [RHRHU08, Question 3.1], they ask about the existence of transitive Anosov diffeomorphisms on non-compact manifolds. And in [BFM25, BBM24], while the axiomatic definition of Anosov-like actions does not require (co)-compactness, compactness of the manifold is used to prove that (pseudo)-Anosov flows gives rise to Anosov-like action, making it natural to ask whether or not compactness is necessary.
@@
Let $M$ be a compact 3-manifold that admits a transitive Anosov flow. If $M$ is not a graph manifold, then there exists an infinite cover of $M$ on which the lifted flow is transitive.
@@ 1.1. Transitivity on abelian covers: A homological condition
In [Sha93], Sharp studied homologically full Anosov flows that can be defined as transitive Anosov flows $\varphi$ on $M$ such that every element of $H_{1}(M;\mathbb{Z})$ is represented by a periodic orbit of $\varphi$. In [DS21], they extended this definition in the following way: Given $Y\to M$ a regular abelian cover of $M$ with associated covering group $G$, then $G/\mathrm{Torsion}$ can be identified with a quotient of $H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$ and they call a transitive Anosov flow on $M$ $Y$-full if every class in $G$ is represented by a periodic orbit of $\varphi$. It is shown in [DS21, Lemma 7.4] that if the lift $\varphi_{Y}\colon Y\to Y$ is transitive then $\varphi$ is $Y$-full.
Gogolev and Rodriguez Hertz proved in [GRH24, Theorem 2.5], the converse of that statement when $Y$ is the maximal abelian cover. We extend their argument to any abelian cover, thus obtaining the following homological condition for transitivity.
@@ 1.2. Transitivity on regular covers: the $\mathbb{R}$-covered case
An Anosov flow $\varphi$ on a $3$-manifold $M$ is called $\mathbb{R}$-covered if the leaf space of its stable, or unstable, foliation is homeomorphic to $\mathbb{R}$ (see Section 2 for details). This class of Anosov flows classically contains all suspension of Anosov diffeomorphisms on the torus, geodesic flows of negatively curved surfaces, as well as many more examples, for instance all Anosov flows preserving a contact structure. For this class of Anosov flows, we have the striking result that their lifts to regular covers are either transitive or all orbits are wandering:
@@
This result uses the recent work of the first author with C. Bonatti and K. Mann in [BBM24] about non-transitive Anosov-like actions. Theorem 4.5 gives another consequence of that work for some $-\mathbb{R}$-covered flows.
@@ 1.3. Transitive Anosov flows and the homotopical characterization of suspension flows
As far as we are aware, there has not been a lot of progress on this question, except for a result showing the non-existence of such examples under additional regularity assumptions; see [OD24].
While we do not answer that question for Anosov diffeomorphisms here, we answer one version for flows, showing that there exists (a lot of) Anosov flows on non-compact surfaces that satisfy what is known to be a homotopical characterization of suspensions in the compact case. More precisely, a result due to work of Barbot and Fenley (see e.g., [BBGRH21, Theorem 2.15] for a proof) gives the following characterization of suspensions.
@@
Let $\varphi$ be an Anosov flow on a compact $3$-manifold $M$. Then $\varphi$ is orbit equivalent [^2] Two flows $\varphi,\psi\colon M\to M$ are orbit equivalent if there is a homeomorphism of $M$ that takes the orbits of $\varphi$ to the orbits of $\psi$. to the suspension of an Anosov diffeomorphism on the torus if and only if no two periodic orbits $\alpha$, $\beta$ of $\varphi$ are freely homotopic (via a non-trivial homotopy if $\alpha=\beta$).
@@
There exists transitive Anosov flows $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ on non-compact manifolds $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ such that every periodic orbit of $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is alone in its free homotopy class.
@@ Acknowledgement.
The authors thanks Katie Mann for many helpful discussions. Both authors were partially supported by the NSERC grants ALLRP 598447 - 24 and RGPIN-2024-04412.
@@ 2. Anosov-like actions and structure of the orbit space
Recall that a (pseudo-)Anosov flow $\varphi$ on a 3-manifold $M$ induces an action of $\pi_{1}(M)$ on the orbit space $P_{\varphi}$ of $\varphi$. The orbit space is defined as the quotient of the universal cover $\widetilde{M}$ of $M$ under the relation “being on the same orbit of the lifted flow $\widetilde{\varphi}$”. A fundamental result of Barbot [Bar95] and Fenley [Fen94] is that the orbit space is a topological plane. The stable and unstable foliations $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$ associated to $\varphi$ lift to foliations ~$\widetilde{\mathcal{F}}^{s}$, ~$\widetilde{\mathcal{F}}^{u}$ of $\widetilde{\varphi}$, which then descend to a pair of transverse foliations of $P_{\varphi}$, giving it a structure of a bifoliated plane $(P_{\varphi},\widetilde{\mathcal{F}}^{s},\widetilde{\mathcal{F}}^{u})$. Moreover, the induced action of $\pi_{1}(M)$ on the orbit space is an example (see [BBM24, Proposition 2.2]) of the general class of Anosov-like action introduced in [BFM25, BBM24].
@@
[…]
@@
Among the axioms in Definition 2.1, (A1) - (A3) reflect the fundamental properties of $\pi_{1}(M)\curvearrowright(P_{\varphi},\widetilde{\mathcal{F}}^{s},\widetilde{\mathcal{F}}^{u})$ such as hyperbolicity and density of the leaves of periodic orbits. However, the proof of [Fen98] that the induced action of a pseudo-Anosov flow satisfy Axiom (A4) relies in an essential way on the compactness of the 3-manifold $M$.
@@
[…]
@@
It is a classical result that the leaf space of a (non-singular) foliation of a plane is a connected, simply-connected, non-compact, and not necessarily Hausdorff 1-manifold; see [CC03]. In the context of Anosov flows, we have the following standard terminology.
@@
    $\mathcal{P}$ is isomorphic [^3] Here, by isomorphic, we mean a foliation-preserving homoeomorphism. to $\mathbb{R}^{2}$ foliated by horizontal and vertical lines, in which case $\mathcal{P}$ is called the trivial plane.
@@
    $\mathcal{P}$ is isomorphic to the strip $\{(x,y)\in\mathbb{R}^{2}\mid x-1<y<x+1\}$ foliated by horizontal and vertical lines, in which case $\mathcal{P}$ is called a skew plane.
@@ 3. Homological conditions for transitivity
In the next two subsections, we will prove Theorem B and deduce (a more general version of) Theorem A.
@@ 3.1. Y-full flows
Let $\varphi\colon M\to M$ be an Anosov flow, and let $Y$ be an abelian regular cover of $M$ with an abelian group of covering transformations $G$. The flow $\varphi$ is said to be Y-full if every class in $G$ is represented by a periodic orbit of $\varphi$.
In the particular case where $Y=M_{\mathrm{ab}}$ is the maximal abelian cover of $M$, i.e., when $G=H_{1}(M;\mathbb{Z})$, then being $M_{\mathrm{ab}}$-full is called being homologically full; see [Sha93]. As we recalled in the introduction, Theorem B is an extension of a result of Gogolev and Rodriguez Hertz in [GRH24] where they proved that a transitive Anosov flow is homologically full if and only if its lift to the maximal abelian cover is transitive. Here we adapt and extend their argument to deal with other abelian covers.
Note that any abelian cover $Y$ is further covered by $M_{\mathrm{ab}}$, so $G/\mathrm{Torsion}$ can always be identified with a quotient of $H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$.
@@
Since $G/\mathrm{Torsion}\cong(H_{1}(M;\mathbb{Z})/\mathrm{Torsion})/K$, where $K$ is the kernel of some surjective map from $H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$ to $G/\mathrm{Torsion}$, we can consider any class $[gK]$ in $G/\mathrm{Torsion}$ as a class $[g]$ in $H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$.
Fix a small $\varepsilon>0$, and let $0<\delta<\varepsilon$ be a constant for which the Anosov closing lemma applies (see e.g., [FH19, Theorem 5.3.10]), i.e., if a an orbit segment of $\varphi$ comes back at distance at most $\delta$ from its starting point, then there exists a periodic orbit that $\varepsilon$-shadows it.
Now let $\gamma$ be an $\delta$-dense periodic orbit of $\varphi$. If $[\gamma]$ is trivial in $G$, then lifts of $\gamma$ to $Y$ are closed orbits and their union form an $\delta$-dense family of close orbits of $\varphi_{Y}$. So for the rest of this proof, suppose that $[\gamma]$ is nontrivial in $G$. By the assumption that $\varphi$ is $Y$-full, we know there exists a periodic orbit $\gamma^{\prime}$ of $\varphi$ such that $[\gamma]+[\gamma^{\prime}]$ is trivial in $G$. Consider $[\gamma]$ and $[\gamma^{\prime}]$ as elements of $H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$.
Since $\gamma$ is $\delta$-dense, for any $y\in\gamma^{\prime}$, there exists some $x\in\gamma$ such that $y$ is contained in the $\delta$-neighborhood of $x$. By the Anosov closing lemma, there exists a periodic orbit $\eta$ of $\varphi$ in the $\varepsilon$-neighborhood of the pseudo-orbit obtained from concatenating $\gamma$ and $\gamma^{\prime}$ through $x$ and $y$. It follows that $\eta$ is $\varepsilon$-dense in $M$. Using Fried’s “pair of pants” construction (see [Fri83, §2]), we can construct an immersed surface whose boundary consists of $\gamma,\gamma^{\prime}$, and $\eta$. Then $\eta$ is homologous to $\gamma+\gamma_{1}$, so $[\eta]=[\gamma]+[\gamma^{\prime}]$ is trivial in $G$. Then $\eta$ lifts to an $\varepsilon$-dense union of closed orbits of $\varphi_{Y}$ in $Y$. This proves that $\varphi_{Y}$ is transitive. ∎
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow. Suppose that there exists linearly independent elements $[\alpha_{1}],\dotsc,[\alpha_{n}]\in H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$ such that for each $i=1,\dots,n$, both $[\alpha_{i}]$ and $-[\alpha_{i}]$ are represented by periodic orbits of $\varphi$. Let $A=\mathrm{span}(\{[\alpha_{1}],\dotsc,[\alpha_{n}]\})$, and write $H_{1}(M;\mathbb{Z})=A\oplus K$ for some normal subgroup $K\lhd H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$. Let $Y$ be the abelian cover of $M$ associated to $K$, i.e., the cover whose group of deck transformations $G\cong(H_{1}(M;\mathbb{Z})/\mathrm{Torsion})/K$. Then the lifted flow $\varphi_{Y}\colon Y\to Y$ is transitive, and in particular, $\varphi$ is $Y$-full.
@@
This theorem tells us that as soon as there exists some nontrivial $[\alpha]\in H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$ such that both $[\alpha]$ and $-[\alpha]$ are represented by some periodic orbits of $\varphi$, then for any infinite cyclic cover $Y$ of $M$ associated to an element $[\delta]\in H^{1}(M;\mathbb{Z})$ with $[\delta]([\alpha])\neq 0$, the lifted flow $\varphi_{Y}$ is transitive.

### v2

@@ Abstract.
In this article we study topological transitivity of Anosov flows on non-compact 3-manifolds. We provide both homological and homotopical conditions under which lifts to infinite covers of transitive Anosov flows stay transitive. In particular, we show that most transitive Anosov flows admit a transitive lift to an infinite cover, and we build a family of examples of transitive Anosov flows on non-compact manifolds that satisfy a homotopical characterization of suspension flows.
2020 Mathematics Subject Classification:
Primary 37D20, 37D05; Secondary 37C27
@@ 1. Introduction
The topological study of Anosov flows, or Anosov diffeomorphisms, has been so far almost exclusively limited to the case where the ambient manifold is compact. Some exceptions to that rule can be found in the works on Anosov diffeomorphisms on the plane or open surfaces, e.g., [GN15, Mat21, Men77, BOD25], as well as recently, in the introduction and study of Anosov-like actions [BFM25, BBM24], where compactness is not explicitly needed.
@@
One of the issues appearing when considering Anosov dynamics in the non-compact setting is that the non-wandering set, i.e., the part of the manifold where the interesting dynamics appear, can be small or even trivial. (A classical example is that there exists a complete metric making the translation on $\mathbb{R}^{2}$ Anosov [Whi73].) So for the dynamics to be interesting, one would like to be able to ensure that the non-wandering set is large enough. In this article, we consider regular infinite covers (see Definition 3.1) of compact manifolds, as they are a very large and natural source of examples of Anosov flows in the non-compact setting and address the following general question:
@@
Let $\varphi$ be a transitive Anosov flow on a compact $3$-manifold $M$. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be any infinite cover of $M$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ the lifted flow. How common is it for $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ to be transitive?
This question of existence and/or abundance of transitive Anosov flows in non-compact manifolds, or infinite covers of compact manifolds has appeared more or less explicitly in several different works. First, in the context of geodesic flows for negatively curved metrics, the question of transitivity (both in the compact and non-compact case), was an historically important one that was resolved in the 1930s (see [Hed39]) in low dimension and, in greater generality and any dimensions, in the early 1970s [Ebe72]. Much more recently, Dougall and Sharp consider in [DS21] certain Abelian covers $Y$ of an Anosov flow $\varphi$ on $M$ and assume the transitivity of the lifted flow $\varphi_{Y}$, making it natural to wonder whether or not this assumption is necessary [^1] Contrarily to this article, the dimension of the manifolds they consider in [DS21] is arbitrary.. In [RHRHU08, Question 3.1], they ask about the existence of transitive Anosov diffeomorphisms on non-compact manifolds. And in [BFM25, BBM24], while the axiomatic definition of Anosov-like actions does not require (co-)compactness, compactness of the manifold is used to prove that (pseudo-)Anosov flows gives rise to Anosov-like action, making it natural to ask whether or not compactness is necessary.
@@
Let $M$ be a compact 3-manifold that admits a transitive Anosov flow. If $M$ is not a graph manifold [^2] A graph manifold is a $P^{2}$-irreducible (irreducible and contains no two-sided $\mathbb{R}\mathrm{P}^{2}$’s) $3$-manifold such that its JSJ decomposition consists only of Seifert fibered pieces, see e.g., [Hat23]., then there exists an infinite cover of $M$ on which the lifted flow is transitive.
@@ 1.1. Transitivity on abelian covers: A homological condition
In [Sha93], Sharp studied homologically full Anosov flows that can be defined as transitive Anosov flows $\varphi$ on $M$ such that every element of $H_{1}(M;\mathbb{Z})$ is represented by a periodic orbit of $\varphi$. In [DS21], they extended this definition in the following way: Given $Y\to M$ a regular abelian cover of $M$ with associated covering group $G$, then $G$ can be identified with a quotient of $H_{1}(M;\mathbb{Z})$, and they call a transitive Anosov flow on $M$ $Y$-full if every class in $G$ is represented by a periodic orbit of $\varphi$. It is shown in [DS21, Lemma 7.4] that if the lift $\varphi_{Y}\colon Y\to Y$ is transitive then $\varphi$ is $Y$-full.
Gogolev and Rodriguez Hertz proved, in [GRH24, Theorem 2.5], the converse of that statement when $Y$ is the maximal abelian cover. We extend their argument to any abelian cover, thus obtaining the following homological condition for transitivity.
@@ 1.2. Transitivity on regular covers: the $\mathbb{R}$-covered case
An Anosov flow $\varphi$ on a $3$-manifold $M$ is called $\mathbb{R}$-covered if the leaf space of its stable, or unstable, foliation is homeomorphic to $\mathbb{R}$ (see Section 2 for details). This class of Anosov flows classically contains all suspension of Anosov diffeomorphisms on the torus, geodesic flows of negatively curved surfaces, as well as many more examples, for instance all Anosov flows preserving a contact structure. For this class of Anosov flows, we have the striking result that their lifts to all regular covers (not just abelian ones) are either transitive or all orbits are wandering:
@@
In particular this theorem extends the classical result of transitivity on regular cover of geodesic flows on negatively curved surfaces [Hed39] to the much wider class of $\mathbb{R}$-covered Anosov flows. Our proof uses the recent work of the first author with C. Bonatti and K. Mann in [BBM24] about non-transitive Anosov-like actions. Theorem 4.5 gives another consequence of that work for some non-$\mathbb{R}$-covered flows.
@@ 1.3. Transitive Anosov flows and the homotopical characterization of suspension flows
As far as we are aware, there has not been a lot of progress on this question, except for a result showing the non-existence of such examples under additional regularity assumptions; see [BOD25].
While we do not answer that question for Anosov diffeomorphisms here, we answer one version for flows, showing that there exists (a lot of) Anosov flows on non-compact manifolds that satisfy what is known to be a homotopical characterization of suspensions in the compact case. More precisely, a result due to the work of Barbot and Fenley (see e.g., [BBGRH21, Theorem 2.15] for a proof) provides the following characterization of suspensions. Before stating it, recall that two flows $\varphi,\psi\colon M\to M$ are said to be orbit equivalent if there is a homeomorphism of $M$ that takes the orbits of $\varphi$ to the orbits of $\psi$.
@@
Let $\varphi$ be an Anosov flow on a compact $3$-manifold $M$. Then $\varphi$ is orbit equivalent to the suspension of an Anosov diffeomorphism on the torus if and only if no two periodic orbits $\alpha$, $\beta$ of $\varphi$ are freely homotopic (via a nontrivial homotopy if $\alpha=\beta$).
@@
There exists transitive Anosov flows $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ on non-compact $3$-manifolds $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ such that every periodic orbit of $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is alone in its free homotopy class.
@@ Acknowledgement.
The authors thank Katie Mann for many helpful discussions. Both authors were partially supported by the NSERC grants ALLRP 598447 - 24 and RGPIN-2024-04412.
@@ 2. Anosov-like actions and structure of the orbit space
Recall that a (pseudo-)Anosov flow $\varphi$ on a 3-manifold $M$ induces an action of $\pi_{1}(M)$ on the orbit space $P_{\varphi}$ of $\varphi$. The orbit space is defined as the quotient of the universal cover $\widetilde{M}$ of $M$ under the relation “being on the same orbit of the lifted flow $\widetilde{\varphi}$”. A fundamental result of Barbot [Bar95] and Fenley [Fen94] is that the orbit space is a topological plane. The stable and unstable foliations $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$ associated to $\varphi$ lift to foliations $\widetilde{\mathcal{F}}^{s}$, $\widetilde{\mathcal{F}}^{u}$ of $\widetilde{\varphi}$, which then descend to a pair of transverse foliations of $P_{\varphi}$, giving it a structure of a bifoliated plane $(P_{\varphi},\widetilde{\mathcal{F}}^{s},\widetilde{\mathcal{F}}^{u})$. Moreover, the induced action of $\pi_{1}(M)$ on the orbit space is an example (see [BBM24, Proposition 2.2]) of the general class of Anosov-like action introduced in [BFM25, BBM24].
@@
A group $G$ acts Anosov-like on a bifoliated plane $\mathcal{P}=(P,\mathcal{F}^{+},\mathcal{F}^{-})$ (possibly with isolated prong singularities) if the action satisfies all of the following axioms:
@@
Among the axioms in Definition 2.1, (A1)(A3) reflect the fundamental properties of $\pi_{1}(M)\curvearrowright(P_{\varphi},\widetilde{\mathcal{F}}^{s},\widetilde{\mathcal{F}}^{u})$ such as hyperbolicity and density of the leaves of periodic orbits. However, the proof of [Fen98] that the induced action of a pseudo-Anosov flow satisfy Axiom (A4) relies in an essential way on the compactness of the 3-manifold $M$.
@@
Suppose a group $G$ acts Anosov-like on a bifoliated plane $\mathcal{P}=(P,\mathcal{F}^{+},\mathcal{F}^{-})$. The following are equivalent:
@@
It is a classical result, originally due to the work of Reeb and Haefliger, that the leaf space of a (non-singular) foliation of a plane is a connected, simply-connected, non-compact, and not necessarily Hausdorff 1-manifold; see e.g., [CC03]. In the context of Anosov flows, we have the following standard terminology.
@@
    $\mathcal{P}$ is isomorphic [^3] Here, by isomorphism we mean a foliation-preserving homeomorphism. to $\mathbb{R}^{2}$ foliated by horizontal and vertical lines, in which case $\mathcal{P}$ is called the trivial plane.
@@
    $\mathcal{P}$ is isomorphic to the strip $\{(x,y)\in\mathbb{R}^{2}\mid x<y<x+1\}$ foliated by horizontal and vertical lines, in which case $\mathcal{P}$ is called a skew plane.
@@ 3. Homological conditions for transitivity
In the next two subsections, we will prove Theorem B and deduce (a more general version of) Theorem A. We start by recalling the notions of covers we are using in this article.
Definition 3.1.
A cover $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ of a manifold $M$ is a quotient $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu=\widetilde{M}/H$ where $H$ is a subgroup of $\pi_{1}(M)$. Let $G=G(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ be the group acting on $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ by deck transformations. The cover is regular if $H$ is a normal subgroup of $\pi_{1}(M)$, in which case we have a natural identification $G\cong\pi_{1}(M)/H$ as well as a base-point independent identification $H\cong\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$. The cover is called abelian if $G$ is an abelian group.
The maximal abelian cover of $M$, denoted $M_{\mathrm{ab}}$, is the regular abelian cover whose fundamental group $\pi_{1}(M_{\mathrm{ab}})\cong[\pi_{1}(M),\pi_{1}(M)]$ is the commutator subgroup of $\pi_{1}(M)$, and the associated group of deck transformations is $H_{1}(M;\mathbb{Z})\cong\pi_{1}(M)/[\pi_{1}(M),\pi_{1}(M)]$.
Note that given any regular abelian cover $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ with associated group of deck transformations $G\cong\pi_{1}(M)/\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$, then the fact that $G$ is abelian implies that $\pi_{1}(M),\pi_{1}(M)]$ must be a subgroup of $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$, and therefore $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ is covered by $M_{\mathrm{ab}}$, hence the name. In particular, the deck transformation group $G$ of such a regular abelian cover can always be identified with a quotient of $H_{1}(M;\mathbb{Z})\cong\pi_{1}(M)/[\pi_{1}(M),\pi_{1}(M)]$.
@@ 3.1. Y-full flows
Let $\varphi\colon M\to M$ be an Anosov flow, and let $Y$ be a regular abelian cover of $M$ with group of deck transformations $G$. The flow $\varphi$ is said to be Y-full if every class in $G$ is represented by a periodic orbit of $\varphi$.
In the particular case where $Y=M_{\mathrm{ab}}$ is the maximal abelian cover of $M$, then being $M_{\mathrm{ab}}$-full is called being homologically full; see [Sha93]. As we recalled in the introduction, Theorem B is an extension of a result of Gogolev and Rodriguez Hertz in [GRH24] where they proved that a transitive Anosov flow is homologically full if and only if its lift to the maximal abelian cover is transitive. Here we adapt and extend their argument to deal with other abelian covers.
To prove Theorem B, it is enough to only consider non-torsion elements in $G$. For simplicity we introduce the following:
Notation.
For a group $G$, we denote by $G_{T}$ the torsion subgroup of $G$, i.e., $G_{T}=\mathrm{Tor}(G)$, and by $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$ the quotient group $G/G_{T}$. We also write $H_{1}(M)$ for $H_{1}(M;\mathbb{Z})$. In particular, $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}=H_{1}(M)/H_{1}(M)_{T}=H_{1}(M;\mathbb{Z})/\mathrm{Tor}(H_{1}(M;\mathbb{Z}))$.
Recall from the beginning of this section that $G$ can always be identified with some quotient of $H_{1}(M)$, thus $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$ is also identified with some quotient of $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$.
@@
Since $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}\cong\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}/K$, where $K$ is the kernel of some surjective map from $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$ to $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$, we can consider any class $[gK]$ in $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$ as a class $[g]$ in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$.
Fix a small $\varepsilon>0$, and let $0<\delta<\varepsilon$ be a constant for which the Anosov closing lemma applies (see e.g., [FH19, Theorem 5.3.10]), i.e., if an orbit segment of $\varphi$ comes back at distance at most $\delta$ from its starting point, then there exists a periodic orbit that $\varepsilon$-shadows it.
Now let $\gamma$ be an $\delta$-dense periodic orbit of $\varphi$. If $\gamma]$ is trivial in $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$, then lifts of (a power of) $\gamma$ to $Y$ are closed orbits and their union form an $\delta$-dense family of close orbits of $\varphi_{Y}$. So for the rest of this proof, suppose that $\gamma]$ is nontrivial in $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$. By the assumption that $\varphi$ is $Y$-full, we know there exists a periodic orbit $\gamma^{\prime}$ of $\varphi$ such that $\gamma]+[\gamma^{\prime}]$ is trivial in $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$. Consider $\gamma]$ and $\gamma^{\prime}]$ as elements of $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$.
Since $\gamma$ is $\delta$-dense, for any $y\in\gamma^{\prime}$, there exists some $x\in\gamma$ such that $y$ is contained in the $\delta$-neighborhood of $x$. By the Anosov closing lemma, there exists a periodic orbit $\eta$ of $\varphi$ in the $\varepsilon$-neighborhood of the pseudo-orbit obtained from concatenating $\gamma$ and $\gamma^{\prime}$ through $x$ and $y$. It follows that $\eta$ is $\varepsilon$-dense in $M$. Using Fried’s “pair of pants” construction (see [Fri83, §2]), we can construct an immersed surface whose boundary consists of $\gamma,\gamma^{\prime}$, and $\eta$. Then $\eta$ is homologous to $\gamma+\gamma_{1}$, so $\eta]=[\gamma]+[\gamma^{\prime}]$ is trivial in $\mkern 1.5mu\overline{\mkern-1.5muG\mkern-1.5mu}\mkern 1.5mu_{T}$. Then $\eta$ lifts to an $\varepsilon$-dense union of closed orbits of $\varphi_{Y}$ in $Y$. This proves that $\varphi_{Y}$ is transitive. ∎
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow. Suppose that there exists linearly independent nontrivial elements $\alpha_{1}],\dotsc,[\alpha_{n}]\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$ such that for each $i=1,\dots,n$, both $\alpha_{i}]$ and $\alpha_{i}]$ are represented by periodic orbits of $\varphi$. Let $A=\mathrm{span}(\{[\alpha_{1}],\dotsc,[\alpha_{n}]\})$, and write $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}=A\oplus K$ for some normal subgroup $K\lhd\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$. Let $Y$ be the abelian cover of $M$ associated to $K$, i.e., the cover whose group of deck transformations $G\cong\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}/K$. Then the lifted flow $\varphi_{Y}\colon Y\to Y$ is transitive, and in particular, $\varphi$ is $Y$-full.
@@
This theorem tells us that as soon as there exists some nontrivial $\alpha]\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$ such that both $\alpha]$ and $\alpha]$ are represented by some periodic orbits of $\varphi$, then for any infinite cyclic cover $Y$ of $M$ associated to an element $\delta]\in H^{1}(M;\mathbb{Z})$ with $\delta]([\alpha])\neq 0$, the lifted flow $\varphi_{Y}$ is transitive.